# Architectural Patterns for Designing Quantum Artificial Intelligence Systems

### v2

@@ Abstract
{highlights}
The unique software architectural solutions in quantum AI primarily focus on determining which tasks to delegate to a quantum computer (the quantum-classic split) and how quantum and classical software components communicate (interface and middleware).
This work identified existing architectural patterns that focus on integrating quantum components into inference engines and organising the interaction between classical components, quantum components, and hardware.
We anticipate that the pattern catalogue developed in this work will serve as guidance for software architects to effectively utilise quantum components in their projects.
Quantum AI systems offer several benefits, such as accelerating the inference and training stages, enhancing robustness against noise in data and adversarial attacks, and achieving better accuracy. However, most of the collected evidence confirming these advantages is obtained from simulations on classical computers.
@@ 2 Research Methodology
S3 A rigorous and robust quantum speed-up in supervised machine learning Liu et al. (2021)
@@
Figure 1 illustrates our evidence-based research methodology for identifying architectural patterns. The methodology is based on a systematic mapping study and is adhering to the guidelines outlined in Petersen et al. (2015). The framework consists of two main steps: pattern identification and pattern classification. The pattern identification step encompasses several substeps, including paper search, paper collection, snowballing, quality assessment, and paper analysis. We specifically target high-quality academic papers related to architectural patterns in quantum computing systems from electronic databases such as the ACM Digital Library, [^1] https://dl.acm.org/ IEEE Xplore, [^2] https://ieeexplore.ieee.org/Xplore/home.jsp SpringerLink, [^3] https://link.springer.com/ Nature Portfolio, [^4] https://www.nature.com/nature-portfolio and Google Scholar. [^5] https://scholar.google.com/ The search string for all databases was composed of two parts, one related to quantum AI and the other to software engineering:
@@
The collected papers were filtered out based on a series of preset criteria (further details are provided in the protocol available in the supplementary materials). We also performed both forward and backward snowballing processes. The quality assessment of a publication’s reliability and relevance to our SMS is conducted by answering the following questions with ’yes’, ’no’, or ’partially’:
@@ 3 Reference architecture, role of architectural patterns in quantum AI and related works
Figure 3: Quantum circuits used in quantum machine learning applications: a) variational quantum circuit, b) variational quantum circuit implementing data re-uploading Schetakis et al. (2022) c) quantum kernel of the support vector machine Liu et al. (2021), d) quantum convolutional neural network Cong et al. (2019), and e) deep quantum neural network Beer et al. (2020, 2021).
@@
Here, we briefly outline the most common examples of quantum circuit architectures that may arise as a result of the quantum software design process for AI systems. Figure 3a represents a common and general design for variational quantum circuits. The input data, represented by the vector $\mathbf{x}$, is encoded into a quantum state via the parameterised unitary operator $U_{\phi}$. There are a variety of possible designs for this unitary operation and corresponding circuitry characterised by different time complexity and robustness. Among them are basis encoding, amplitude encoding, angle encoding, and IQP-style encoding Weigold et al. (2021a); LaRose and Coyle (2020); Rath and Date (2023). The prepared quantum state is then modified by a circuit implementing another unitary operation $V(\boldsymbol{\theta})$, which contains learnable parameters $\boldsymbol{\theta}$. The unitary operations $U_{\phi}$ and $V(\boldsymbol{\theta})$ consist of rotational single qubit gates and multiqubit controlled gates that introduce quantum entanglement Nielsen and Chuang (2010). The learnable parameters $\boldsymbol{\theta}$ are usually rotation angles in the single-qubit rotational gates. The output is obtained by measuring the quantum states, which converts the quantum states back into classical data. Data encoding does not need to be implemented just once. One of the design solutions employed alongside the variational quantum circuit is a so-called data re-uploading Schetakis et al. (2022). Data re-uploading is a subclass of quantum embedding implemented repeatedly in several quantum layers, as is illustrated in Fig. 3b. This approach can enhance both the expressivity and trainability of the model. The circuit in Fig. 3c presents a quantum algorithm that computes the kernel function for the implementation of a quantum-kernel support vector machine. It computes a feature map based on the distance between two data points, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. The circuit in Fig. 3d illustrates the design of a quantum convolutional neural network as proposed in Cong et al. (2019). This neural network includes analogs of classical pooling layers, which are implemented using the unitary gates $V_{1}$. These gates have parameters that are determined by measurements from other qubits. The circuit in Fig. 3d represents the design of a quantum deep neural network Beer et al. (2020, 2021). In this design, each layer of the quantum neural network, depicted from left to right in the inset in Fig. 3e, corresponds to a multi-qubit channel arranged from top to bottom in the circuit diagram.
@@ 4 Pattern catalog for quantum AI systems (RQ1)
This section provides a detailed description of the architectural patterns identified during the systematic mapping study. Each pattern describes a high-level solution to a challenge encountered when designing, implementing, and/or managing AI systems Lu et al. (2023a). In this work, similar to Lu et al. (2024a), each pattern is formally represented using a template that includes the following components: name (with an optional graphical representation), summary, description of the associated problem, solution provided by the pattern, benefits, drawbacks, and known uses. In this work, we classify architectural patterns specific to quantum AI systems into two categories: quantum-classical split patterns (subsection 4.1) and quantum middleware patterns (subsection 4.2) .
@@ 4.1 Quantum-classic split architectural patterns (SP)
Problem: The problem to solve is enhancing expressiveness of the quantum-enhanced model and increasing the number of trainable parameters within the constraints imposed by NISQ-era quantum computers, which limit the number of qubits and the depth of quantum circuits.
Solution: Given the constraints of NISQ computers, the expressibility of the model and inference accuracy can be enhanced by integrating several quantum components, which are interconnected by classical channels. This setup functions similarly to how additional layers of neurons in deep artificial neural networks improve performance by allowing for more complex and hierarchical feature extraction. In particular implementations of this pattern, multiple quantum components can function as individual quantum neurons [S6], layers of neurons [S17, S35, S73, S110], or they can perform various tasks in the inference pipeline, such as feature extraction and subsequent feature classification in quantum support vector machines [S111].
@@
Solution: The limitations of NISQ-era quantum hardware, particularly the restricted depth of quantum circuits, can be mitigated by embedding a quantum inference engine into a pipeline with classical inference engines, replacing several intermediate layers of the inference pipeline. In this approach, the classical components reduce diminsionality input data, performing initial stages of the feature extraction, and implement classification of the refined features, but also processes data output from the quantum component interpreting results of the quantum state measurements. This solution can be viewed as a combination of the Quantum Head pattern and Quantum Feature Engineering where classical components participate both in quantum embedding for feature extraction and feature processing. This pattern is often utilised in applications addressing classification problems. In a classifier, classical components reduce the dimensionality of the input data through feature extraction and interpret the output from the quantum component [S4; S16; S24; S29; S33-34; S36; S39; S49; S95; S112]. Additionally, the quantum intermediate layer can serves as a blueprint for systems implementing quantum variational autoencoders [S28, S30].
@@ 5 Rationale for using quantum computing in AI systems (RQ2)
Inference speed-up. Many papers report the potential for algorithmic inference speed-up in quantum AI through the use of quantum neural networks [S26; S77; S113; S125] and quantum kernels for support vector machines [S3; S98; S111]. This speed-up is typically associated with the reduced time complexity reported for quantum algorithms compared to their classical analogs [S26; S77; S98; S111; S113]. Note that, given the slow operational speed of existing quantum gates, transferring the speedup associated with reduced time complexity in algorithms to actual speedup on quantum hardware, relative to classical systems, may require problem sizes that are unrealistically large. Several numerical experiments on classical simulations of quantum computers confirms faster execution times compared to classical algorithms for quantum neural networks [S36] and quantum kernels [S56]. The fact that speed-up can be achieved even in simulations lays the foundation for quantum-inspired machine learning systems. Quantum-enhanced kernels for support vector machines have been also successfully run on an IBM quantum computer, although their execution times were not reported [S98; S111].
Reduce model parameters. Another advantage that gate-based quantum computing systems offer to AI is a substantial reduction in the number of model parameters without loss of performance [S10-11; S16; S23; S26; S37-38; S63; S65; S87; S112]. Due to superposition and entanglement, quantum models can be more expressive with fewer parameters. This reduction in parameters can lead to speed-ups during both the training and inference stages.
@@
Quantum sampling supremacy. Quantum mechanical systems can produce probability distributions that exhibit quantum correlations, which are challenging to capture using classical models [S22]. Quantum generative models based on quantum neural networks facilitates efficient learning and loading of generic probability distributions with an exponential speed-up compared to classical algorithms [S22; S71; S100]. While this capability may not have immediate practical applications at the model level, it serves as a compelling demonstration of “quantum learning supremacy,” which holds significant interest for the advancement of quantum AI [S71; S100].
@@ 6 Discussion
The number of publications corresponding to each of the architectural patterns described above is shown in Fig. 11. The most widely used pattern is the quantum monolith. This architecture is the simplest and is well-suited for demonstrating proof-of-concepts and conducting experiments in the field of QML. The next most used patterns are quantum head and intermediate quantum layer. These architectural solutions are closely tied to real-world AI applications for NISQ-era quantum computers. They are well-suited for high-dimensional data and offer promising advantages in the near future.
Uploading large amounts of data directly into quantum computers poses challenges for NISQ-era systems due to limitations in the number of available qubits and their coherence time. As a result, the quantum feature extraction pattern is less commonly used. This pattern is mostly utilised in quantum support vector machines, where the quantum component operates directly on data to compute the feature map. The limitations related to a small number of qubits are somewhat alleviated by quantum convolutional filters (Quanvolution), where the filter scans over the input data layer, processing only a small amount of data at a time. These systems are expected to significantly reduce the number of learnable parameters without loss of expressibility. However, this potential has not yet been fully realised on NISQ computers since it requires deep integration and communication between classical and quantum components.
@@
Distributed quantum computing Distributed quantum computing presents an architecturally scalable solution for building large-scale quantum computing systems, addressing the limitations of NISQ systems. Several experiments have been conducted in this direction. Several authors Wu et al. (2023); Kawase (2024) explored the distribution of partitioned features across multiple spatially separated quantum circuits (nodes) and reported a slight decrease in performance compared to a monolithic quantum neural network. While dataset distribution shares similarities with classical approaches, the distribution of quantum models is distinct due to the need for distributed quantum networks to establish entanglement across different nodes Pira and Ferrie (2023). The inter-node two-qubit gates can be managed using three approaches: terminating two-qubit gates with measurements, approximating them with local gates, or establishing a coherent quantum communication channel between non-local qubits Pira and Ferrie (2023).
Physically, the latter approach involves long-range communication with fibre optics to establish entangled Einstein-Podolsky-Rosen photon pairs Häner et al. (2021); Wu et al. (2023). Communication between nodes cannot be broadband; thus, optimal partitioning of quantum circuits to minimise inter-node two-qubit gates is essential. This optimisation can be automated at the compiler level, as implemented in the QuComm framework Wu et al. (2023), or manually using the Quantum Message Passing Interface, analogous to MPI functions in multi-processing programming Häner et al. (2021). Additionally, the concept of federated learning, which has gained significant attention in classical deep learning, is considered beneficial for quantum machine learning McMahan et al. (2017); Li et al. (2021c); Larasati et al. (2022); Xia and Li (2021); Chehimi and Saad (2022).
@@ 7 Threats to validity
Many papers on quantum software report results from experiments conducted on simulators rather than actual quantum computers. There is limited literature on QML in the context of hardware noise or errors, as well as on QML architectures incorporating quantum error correction. Additionally, most of the works considered in this review pertain to prototypes and experimental software systems that have not yet reached the production stage. The lack of real-world production experience in quantum software could influence existing architectural patterns in two ways: some patterns may not be adopted in production, while new patterns may emerge. However, the theoretical justification for the usefulness of the identified patterns is strong enough to suggest their applicability in the production stage as well.
@@
Another challenge discussed in Ronnow et al. (2014) concerns the experimental evaluation of quantum speedup as reported in the literature. The authors argue that, to properly distinguish parallel speedup from quantum speedup, it is crucial to scale hardware resources identically for both quantum and classical devices during comparisons. Moreover, it is important to avoid conflating inefficiencies at small problem sizes with signs of quantum speedup.

### v3

@@ 2 Research Methodology
S3 A rigorous and robust quantum speed-up in supervised machine learning Liu et al. (2021b)
@@
Figure 1 illustrates our evidence-based research methodology for identifying architectural patterns. The methodology is based on a systematic mapping study and is adhering to the guidelines outlined in Petersen et al. (2015). The framework consists of two main steps: pattern identification and pattern classification. The pattern identification step encompasses several substeps, including paper search, paper collection, snowballing, quality assessment, and paper analysis. We specifically target high-quality academic papers related to architectural patterns in quantum computing systems from electronic databases such as the ACM Digital Library, [^1] https://dl.acm.org/ IEEE Xplore, [^2] https://ieeexplore.ieee.org/Xplore/home.jsp SpringerLink, [^3] https://link.springer.com/ Nature Portfolio, [^4] https://www.nature.com/nature-portfolio and Google Scholar. [^5] https://scholar.google.com/ Broad-scoped scientific databases like GoogleScholar and Scopus can be used interchangeably. While Scopus serves as a general-purpose database in our research methodology, it often retrieves duplicate records, particularly when articles are published in multiple versions or indexed under slightly different author names or journal titles Liu et al. (2021a). This can lead to inefficiencies in the review process, which we aim to minimize. For this reason, we use GoogleScholar to ensure comprehensive coverage while maintaining methodological efficiency. The search string for all databases was composed of two parts, one related to quantum AI and the other to software engineering:
@@
The collected papers were filtered out based on a series of preset criteria (further details are provided in the protocol available in the supplementary materials). Backward snowballing was employed during the quality assessment phase. In this phase, we analysed references from the primary sources, screened them based on title and context within the citing papers, collected relevant studies, and added them to the list for further quality assessment. Forward snowballing was applied to high-impact papers. We examined the publications citing these sources, screened them by title and abstract, collected relevant studies, and incorporated them into the list for quality assessment. The quality assessment of a publication’s reliability and relevance is conducted by answering the following questions with ’yes’, ’no’, or ’partially’:
@@
The sources listed in Table 1) are relevant to research questions RQ1 and RQ2 and are analysed in Sections 4 and 5, respectively, to address them.
@@ 3 Reference architecture, role of architectural patterns in quantum AI and related works
Figure 3: Quantum circuits used in quantum machine learning applications: a) variational quantum circuit, b) variational quantum circuit implementing data re-uploading Schetakis et al. (2022) c) quantum kernel of the support vector machine Liu et al. (2021b), d) quantum convolutional neural network Cong et al. (2019), and e) deep quantum neural network Beer et al. (2020, 2021).
@@
Here, we briefly outline the most common examples of quantum circuit architectures that may arise as a result of the quantum software design process for AI systems. Figure 3a represents a common and general design for variational quantum circuits. A quantum circuit is one of the models of quantum computing serving as a representation of a quantum program. The circuit describes transformations of quantum states of qubits, which are depicted as horizontal lines in Figure 3. At the beginning of the program, all qubit states are typically initialised to a ket vector $|0\rangle$, which represents a vector, expressed using Dirac notation, in a complex vector space Nielsen and Chuang (2010). The input data, represented by the vector $\mathbf{x}$, is encoded into a quantum state via the parameterised unitary operator $U_{\phi}$. There are a variety of possible designs for this unitary operation and corresponding circuitry characterised by different time complexity and robustness. Among them are basis encoding, amplitude encoding, angle encoding, and IQP-style encoding Weigold et al. (2021a); LaRose and Coyle (2020); Rath and Date (2023). The prepared quantum state is then modified by a circuit implementing another unitary operation $V(\boldsymbol{\theta})$, which contains learnable parameters $\boldsymbol{\theta}$. The unitary operations $U_{\phi}$ and $V(\boldsymbol{\theta})$ consist of rotational single qubit gates and multiqubit controlled gates that introduce quantum entanglement Nielsen and Chuang (2010). The learnable parameters $\boldsymbol{\theta}$ are usually rotation angles in the single-qubit rotational gates. The output is obtained by measuring the quantum states, which converts the quantum states back into classical data. Data encoding does not need to be implemented just once. One of the design solutions employed alongside the variational quantum circuit is a so-called data re-uploading Schetakis et al. (2022). Data re-uploading is a subclass of quantum embedding implemented repeatedly in several quantum layers, as is illustrated in Fig. 3b. This approach can enhance both the expressivity and trainability of the model. The circuit in Fig. 3c presents a quantum algorithm that computes the kernel function for the implementation of a quantum-kernel support vector machine. It computes a feature map based on the distance between two data points, $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$. The circuit in Fig. 3d illustrates the design of a quantum convolutional neural network as proposed in Cong et al. (2019). This neural network includes analogs of classical pooling layers, which are implemented using the unitary gates $V_{1}$. These gates have parameters that are determined by measurements from other qubits. The circuit in Fig. 3d represents the design of a quantum deep neural network Beer et al. (2020, 2021). In this design, each layer of the quantum neural network, depicted from left to right in the inset in Fig. 3e, corresponds to a multi-qubit channel arranged from top to bottom in the circuit diagram.
@@ 4 Pattern catalog for quantum AI systems (RQ1)
This section provides a detailed overview of the architectural patterns identified during the systematic mapping study. Each pattern presents a high-level solution to a challenge encountered in the design, implementation, and management of AI systems Lu et al. (2023a). In this work, similar to Lu et al. (2024a), each pattern is formally represented using a template that includes the following components: name (with an optional graphical representation), summary, description of the associated problem, solution provided by the pattern, benefits, drawbacks, and known uses. In this work, we classify architectural patterns specific to quantum AI systems into two categories: quantum-classical split patterns (discussed in a subsection 4.1) and quantum middleware patterns (presented in a subsection 4.2).
@@ 4.1 Quantum-classic split architectural patterns (SP)
Problem: The problem to solve is enhancing expressivity of the quantum-enhanced model and increasing the number of trainable parameters within the constraints imposed by NISQ-era quantum computers, which limit the number of qubits and the depth of quantum circuits.
Solution: Given the constraints of NISQ computers, the expressivity of the model and inference accuracy can be enhanced by integrating several quantum components, which are interconnected by classical channels. This setup functions similarly to how additional layers of neurons in deep artificial neural networks improve performance by allowing for more complex and hierarchical feature extraction. In particular implementations of this pattern, multiple quantum components can function as individual quantum neurons [S6], layers of neurons [S17, S35, S73, S110], or they can perform various tasks in the inference pipeline, such as feature extraction and subsequent feature classification in quantum support vector machines [S111].
@@
Solution: The limitations of NISQ-era quantum hardware, particularly the restricted depth of quantum circuits, can be mitigated by embedding a quantum inference engine into a pipeline with classical inference engines, replacing several intermediate layers of the inference pipeline. In this approach, the classical components reduce diminsionality input data, performing initial stages of the feature extraction, and also processes data output from the quantum component interpreting results of the quantum state measurements. This solution can be viewed as a combination of the Quantum Head pattern and Quantum Feature Engineering where classical components participate both in quantum embedding for feature extraction and feature processing. This pattern is often utilised in applications addressing classification problems. In a classifier, classical components reduce the dimensionality of the input data through feature extraction and interpret the output from the quantum component [S4; S16; S24; S29; S33-34; S36; S39; S49; S95; S112]. Additionally, the quantum intermediate layer can serves as a blueprint for systems implementing quantum variational autoencoders [S28, S30].
@@ 5 Rationale for using quantum computing in AI systems (RQ2)
Inference speed-up. Many papers report the potential for algorithmic inference speed-up in quantum AI through the use of quantum neural networks [S26; S77; S113; S125] and quantum kernels for support vector machines [S3; S98; S111]. This speed-up is typically associated with the reduced time complexity reported for quantum algorithms compared to their classical analogs [S26; S77; S98; S111; S113]. Due to the limited operational speed of current quantum gates, achieving actual performance gains on quantum hardware from the speedup associated with reduced time complexity in algorithms may require impractically large problem sizes. Several numerical experiments on classical simulations of quantum computers confirms faster execution times compared to classical algorithms for quantum neural networks [S36] and quantum kernels [S56]. The fact that speed-up can be achieved even in simulations lays the foundation for quantum-inspired machine learning systems. Quantum-enhanced kernels for support vector machines have been also successfully run on an IBM quantum computer, although their execution times were not reported [S98; S111].
One of the key challenges in achieving practical speedups in QML is the time/gate complexity of encoding classical information into quantum states. A potential solution to this challenge is the use of Quantum Random Access Memory Giovannetti et al. (2008), which can reduce this complexity and improve data access speed. However, quantum memories require fault-tolerant quantum computers and remain an active area of research, with practical, large-scale implementations still yet to be realised. Another approach aims to leverage approximate encoding methods and incorporates variational, genetic, and matrix product state algorithms West et al. (2024).
Reduce model parameters. Another advantage that gate-based quantum computing systems offer to AI is a substantial reduction in the number of model parameters without loss of performance [S10-11; S16; S23; S26; S37-38; S63; S65; S87; S112]. It is anticipated that quantum models with fewer parameters can achieve the same performance as classical models.
@@
Quantum sampling supremacy. Quantum mechanical systems can produce probability distributions that exhibit quantum correlations, which are challenging to capture using classical models [S22]. Quantum generative models based on quantum neural networks facilitates efficient learning and sampling of generic probability distributions with an exponential speed-up compared to classical algorithms [S22; S71; S100]. While this capability may not have immediate practical applications at the model level, it serves as a compelling demonstration of “quantum learning supremacy,” which holds significant interest for the advancement of quantum AI [S71; S100].
@@ 6 Discussion
The number of publications corresponding to each of the architectural patterns described above is shown in Figure 11. The most widely used pattern is the quantum monolith. This architecture is the simplest and is well-suited for demonstrating proof-of-concepts and conducting experiments in the field of QML. The next most used patterns are quantum head and intermediate quantum layer. These architectural solutions are closely tied to real-world AI applications for NISQ-era quantum computers. They are well-suited for high-dimensional data and offer promising advantages in the near future.
Uploading large amounts of data directly into quantum computers poses challenges for NISQ-era systems due to limitations in the number of available qubits and their coherence time. As a result, the quantum feature extraction pattern is less commonly used. This pattern is mostly utilised in the quantum kernel methods, where the quantum component operates directly on data to compute the feature map. In Gil-Fuster et al. (2024a), the universality – and thus expressivity – of embedding quantum kernels has been formally proven. The limitations related to a small number of qubits are somewhat alleviated by quantum convolutional filters, where the filter scans over the input data layer, processing only a small amount of data at a time. These systems are expected to significantly reduce the number of learnable parameters without loss of expressibility. However, this potential has not yet been fully realised on NISQ computers since it requires deep integration and communication between classical and quantum components.
@@
Note that existing alternatives to the quantum monolith architecture are primarily motivated by the limitations of the NISQ era, such as the limited number of quantum resources and noise. However, there are other fundamental challenges that may prevent the quantum monolith from being the optimal architecture, even when fault-tolerant, large-scale quantum computers become a reality. One such challenge is the trade-off between the expressivity and trainability of quantum neural networks, driven by the barren plateau problem. This issue arises due to the extremely high dimensionality of the underlying Hilbert space of quantum states. A more expressive quantum models also require a higher measurement cost in gradient estimation Chinzei et al. (2024). While this dimensionality and expressivity can be reduced through specific quantum circuit designs, such as quantum convolutional neural networks (see Figure 3d), circuits that avoid barren plateaus might be classically simulable, thereby negating the quantum advantage – which is the primary motivation for the quantum monolith architecture Cerezo et al. (2024); Gil-Fuster et al. (2024b).
@@
Distributed quantum computing Distributed quantum computing Caleffi et al. (2024) presents an architecturally scalable solution for building large-scale quantum computing systems, addressing the limitations of NISQ systems. Several experiments have been conducted in this direction. Several authors Wu et al. (2023); Kawase (2024) explored the distribution of partitioned features across multiple spatially separated quantum circuits (nodes) and reported a slight decrease in performance compared to a monolithic quantum neural network. While dataset distribution shares similarities with classical approaches, the distribution of quantum models is distinct due to the need for distributed quantum networks to establish entanglement across different nodes Pira and Ferrie (2023). The inter-node two-qubit gates can be managed using three approaches: terminating two-qubit gates with measurements, approximating them with local gates, or establishing a coherent quantum communication channel between non-local qubits Pira and Ferrie (2023).
Physically, the latter approach involves long-range communication with fibre optics to establish entangled Einstein-Podolsky-Rosen photon pairs Häner et al. (2021); Wu et al. (2023). Communication between nodes cannot be broadband; thus, optimal partitioning of quantum circuits to minimise inter-node two-qubit gates is essential. This optimisation can be automated at the compiler level, as implemented in the QuComm framework Wu et al. (2023), or manually using the Quantum Message Passing Interface, analogous to MPI functions in multi-processing programming Häner et al. (2021).
Additionally, the concept of federated learning, which has gained significant attention in classical deep learning, is considered beneficial for quantum machine learning McMahan et al. (2017); Li et al. (2021c); Larasati et al. (2022); Xia and Li (2021); Chehimi and Saad (2022). It has been shown in Heredge et al. (2024) that QML with federated learning demonstrates the potential for increased privacy, unattainable in classical settings.
@@ 7 Threats to validity
Many papers on quantum software report results from experiments conducted on simulators rather than actual quantum computers. There is limited literature on QML in the context of hardware noise or errors, as well as on QML architectures incorporating quantum error correction. Additionally, most of the works considered in this review pertain to prototypes and experimental software systems that have not yet reached the production stage. The lack of real-world production experience in quantum software could influence existing architectural patterns in two ways: some patterns may not be adopted in production, while new patterns may emerge. There are only a few works that have conducted quantum validation on hardware, and the number of such studies is insufficient for a systematic mapping study. Nevertheless, the papers selected for this study offer valuable insights through theoretical reasoning, simulations, small-scale experiments, and toy models. These contributions allow us to identify potential architectural patterns for quantum AI systems even in the absence of full quantum validation. The theoretical justification for the usefulness of the identified patterns is strong enough to suggest their applicability in the production stage as well.
@@
Another challenge discussed by Ronnow et al. (2014) concerns the experimental evaluation of quantum speedup as reported in the literature. The authors argue that, to properly distinguish parallel speedup from quantum speedup, it is crucial to scale hardware resources identically for both quantum and classical devices during comparisons. Moreover, they emphasize the need to avoid confusing inefficiencies at small problem sizes with true indications of quantum speedup.